Rhic heavy ion run 11
This presentation is the property of its rightful owner.
Sponsored Links
1 / 144

RHIC Heavy Ion Run-11 PowerPoint PPT Presentation


  • 45 Views
  • Uploaded on
  • Presentation posted in: General

RHIC Heavy Ion Run-11. Status Meeting Updates. Heavy Ion Run Final Update 6/24-27. Physics running, 41 stores Tandem intensity issues Access controls failure Access for RF trip. Outstanding items. Next 48 hours 6/27-29. Physics ends 2200 Tuesday 6/28

Download Presentation

RHIC Heavy Ion Run-11

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Rhic heavy ion run 11

RHIC Heavy Ion Run-11

Status Meeting Updates


Heavy ion run final update 6 24 27

Heavy Ion Run Final Update 6/24-27

  • Physics running, 41stores

    • Tandem intensity issues

    • Access controls failure

    • Access for RF trip

Outstanding items


Next 48 hours 6 27 29

Next 48 hours 6/27-29

  • Physics ends 2200 Tuesday 6/28

  • APEX and Power Supply tests until 0600 Wednesday 6/29


Heavy ion run update 6 22 6 24

Heavy Ion Run Update 6/22-6/24

  • Physics running, 28 stores

    • No failed ramps

    • 4 aborted stores

      • ODH HW failure

      • Node card failure

      • Abort kicker trip

      • Power dip

  • Blue injection retuned

  • Reduced store length to 1 hour

  • Conducted vernier scans

Outstanding items

  • Yellow end-ramp efficiency

    • Better overnight

    • Injection damper issues?


Weekend outlook 6 24 27

Weekend outlook 6/24-27

  • Network recovery

  • Continue Physics

  • APEX and MD next week?

  • Last meeting: Monday, noon


Heavy ion run update 6 20 6 22

Heavy Ion Run Update 6/20-6/22

  • Setup started ~1730 Monday

  • Physics started ~1730 Tuesday

    • 6 stores

  • Injection & ramp collimation implemented

  • Correcting injection drift

  • Tandem reverted to optimal Au configuration

Outstanding items

  • Blue injection efficiency

  • Yellow ramp efficiency

  • Injection damper issues?


Next 48 hours 6 22 24

Next 48 hours 6/22-24

  • Continue Physics

    • Store length? Optimal ~1 hour.

    • Can detectors remain on?

  • Status meetings M-W-F


Heavy ion run update 6 17 6 20

Heavy Ion Run Update 6/17-6/20

  • Physics running, 12 stores

    • 2 failed ramps

      • Blue abort prefire

    • 1 aborted store

      • power dip, QLI

  • Tandem foil work conducted behind store Friday.

  • One additional store for AnDY (16081)

  • 200 GeV Physics ended 0800 Monday.

Outstanding items

  • Blue injection efficiency

  • Storage cavity voltage calibration

    • Blue cavities

    • Requires working Blue longitudinal Schottky


27 gev startup 6 20 22

27 GeV startup 6/20-22

  • Starts Monday 1600

    • IR access during hysteresis

    • Beam development follows

      • Re-establish injected beam

      • Ramp with feedback

      • Store setup

  • Status meetings M-W-F


Heavy ion run update 6 15 6 17

Heavy Ion Run Update 6/15-6/17

  • Physics running, 6 stores

    • 1 failed ramp

      • LLRF test ramp

    • 1 aborted store

      • False BLM trip, 7W

  • IP2 crossing angle removed after maintenance

    • Corrected with feedbacks

  • Machine Development: 9 hours

    • 7 hours LLRF

      • Punctuated with 2 HLRF failures/accesses and Blue injection kicker failure

      • Second ramp needed

    • 2 hours AnDY collision setup

Outstanding items

  • Tandem low on stripping foils

  • Blue injection efficiency

  • Blue transition

    • Landau problem again, fill 16071

  • Storage cavity voltage calibration

    • Blue cavities

    • Requires working Blue longitudinal Schottky


Zdc accidental correction

ZDC Accidental correction

  • Number of buckets: Blue filled bunches

  • STAR

    • ZDC Singles East, ch 1

    • ZDC Singles West, ch 2

    • ZDC, ch 3

      Corrected=ZDC-{East*West/[(Blue-Blue/14)*78000]}

  • PHENIX

    • ZDC Singles North, ch 1.1

    • ZDC Singles South, ch 1.2

    • ZDC, ch 16

      Corrected=ZDC-[North*South/(Blue*78000)]


Final 200 gev weekend 6 17 20

Final 200 GeV weekend 6/17-20

  • Continue Physics

    • One more store with AnDY

    • One more set of vernier scans

    • One more test of β* squeeze

  • Tandem foil work behind store

  • APEX Monday 0800-1600

  • √SNN = 27 GeV beam development starts Monday 1600

  • Status meetings M-W-F


Heavy ion run update 6 13 6 15

Heavy Ion Run Update 6/13-6/15

  • Physics running, 10 stores

    • 3 failed ramps

      • Yellow feedback target tunes too close?

      • Yellow debunching before accramp

      • Blue abort kicker prefire + quench

    • 4 aborted stores

      • Yo1-sxf1 trip

      • Yellow abort kicker prefire + quench

      • Wrong Chrom values left from MD

      • Cryo lead flow interlock

  • Beta squeeze, DX ramp test for AnDY work

    • Partially behind Booster RF failure

  • Some Tandem intensity issues

Outstanding items

  • Blue injection efficiency

  • Storage cavity voltage calibration

    • Blue cavities

    • Requires working Blue longitudinal Schottky


Next 48 hours 6 13 15

Next 48 hours 6/13-15

  • LLRF firmware upgrade

    • Dump at 0900 tomorrow to begin

  • AnDY ZDC calibration

    • 2 stores following LLRF work

  • Continue Physics

  • √SNN = 27 GeV beam development next Monday

  • Status meetings M-W-F


Heavy ion run update 6 10 6 13

Heavy Ion Run Update 6/10-6/13

  • Physics running, 14 stores

    • 2 failed ramps

      • Network switch failure, 11A

      • BBQ lock windows not set for ramp

    • 1 early dump

      • Rebucketing failure, HLRF misconfiguration after access

Outstanding items

  • Blue longitudinal problems on ramp

  • Blue injection efficiency

  • Storage cavity voltage calibration

    • Blue cavities

  • LLRF firmware upgrade

    • 4 hrs, week of 6/13

  • AnDY ZDC calibration


27 gev medium energy run notes

27 GeV medium energy run notes

  • Ramp name Au11v6 (work in progress)

    • 13.5 GeV beam energy

    • 3m β* at IP6 and IP8

    • Below transition

  • No rebucketing

  • No longitudinal cooling

  • AP shifts start 0800 Monday 6/20


Next 48 hours 6 13 151

Next 48 hours 6/13-15

  • End of store MD: β* squeeze at IP6

    • Slow ramp to try and eliminate rise in beam decay

  • Continue Physics

    • Continue fine tuning cooling for best performance.

  • √SNN = 27 GeV beam development next Monday

  • Status meetings M-W-F


Heavy ion run update 6 8 6 10

Heavy Ion Run Update 6/8-6/10

  • Physics running, 4 stores

    • no failed ramps

    • 2 early dumps

      • Sextupole temperature problems in 1003B

  • APEX, 16 hours

    • Lost 4 hours to weather and related issues.

Outstanding items

  • Blue longitudinal problems on ramp

    • Although, none on the last few ramps

  • Blue injection efficiency

  • Storage cavity voltage calibration

    • Blue cavities

  • LLRF firmware upgrade

    • 4 hrs, week of 6/13

  • AnDYZDC calibration


Weekend outlook 6 10 13

Weekend outlook 6/10-13

  • Access for PHENIX RPC issues.

  • Continue Physics

    • Continue fine tuning cooling for best performance.

  • NOTE plan to run at √SNN = 27 GeV

  • Status meetings M-W-F


Heavy ion run update 6 6 6 8

Heavy Ion Run Update 6/6-6/8

  • Physics running, 9 stores

    • 1 failed ramp

      • Early ramp loss, LLRF problem

    • 2 aborted stores

      • Yellow abort kicker

      • Yellow RF permit failure

  • Machine development

    • Beta* squeeze test, “MAD-X knob”

    • Some complications to measured results, will try again

Outstanding items

  • Blue longitudinal problems on ramp

  • Blue injection efficiency

  • Storage cavity voltage calibration

    • Blue cavities

  • LLRF firmware upgrade

    • 4 hrs, week of 6/13


Rhic heavy ion run 11

Cooling on vs. off


Next 48 hours 6 6 8

Next 48 hours 6/6-8

  • Continue Physics

    • Continue fine tuning cooling for best performance.

  • Test β* squeeze at end of store.

  • APEX Thursday

    • Note NSRL/Tandem Si beam conflict

  • Status meetings M-W-F


Heavy ion run update 6 3 6 6

Heavy Ion Run Update 6/3-6/6

  • Physics running, 14 stores

    • 3 failed ramps

      • Sextupole trip late in ramp

      • Early Blue loss, RF or node card?

    • 1 aborted store, QD problem

    • Landau cavity issues at transition

  • Machine development

    • Beta* squeeze test, “MAD-X knob”

Outstanding items

  • Blue injection efficiency

  • Storage cavity voltage calibration

    • Blue cavities

  • LLRF firmware upgrade

    • 4 hrs, week of 6/13

  • Stochastic cooling issues

    • BL kicker drive


Next 48 hours 6 6 81

Next 48 hours 6/6-8

  • Continue Physics

    • Continue fine tuning cooling for best performance.

  • APEX Wednesday

    • Note NSRL/Tandem Si beam conflict

  • Status meetings M-W-F


Heavy ion run update 5 27 6 1

Heavy Ion Run Update 5/27-6/1

  • Maintenance, vacuum pump down complete Thursday 0330.

    • STAR, PHENIX polarity reversal

  • Physics running, 5 stores

    • 1 early dump

      • Low rates, Blue vacuum

    • 3 aborted stores

      • Yellow quench after rebucketing, YS5 cavity problem

      • Yellow abort late in store, related to activate?

      • Cryo lead flow

  • Machine development

    • Beta* squeeze test, all planes both IRs

    • Working point scan

    • Yellow RF cavity voltage calibration

Outstanding items

  • Blue injection efficiency

  • Storage cavity voltage calibration

    • Blue cavities

  • LLRF firmware upgrade

    • 4 hrs, week of 6/13

  • Stochastic cooling issues

    • BL kicker drive repair

    • BV not yet operational


Rhic heavy ion run 11

New RF voltage calibrations


Rhic heavy ion run 11

Run-10 Yellow working point

Present Yellow working point


Weekend outlook 6 3 6

Weekend outlook 6/3-6

  • Restore stochastic cooling systems

    • Requires additional access

    • PHENIX access concurrently

  • Continue Physics

    • Vernier scans Friday: early in store

  • Status meetings M-W-F


Heavy ion run update 5 27 6 11

Heavy Ion Run Update 5/27-6/1

  • Physics running, 19 stores

    • 5 failed ramps

      • Early loss attributed to y-qtrim?

      • Yellow RF trip on ramp

      • Fast Blue loss ~10 sec after γt (2x)

      • Unexplained fast Blue loss late in ramp

    • 5 aborted stores

      • PASS failure

      • Blue abort kicker

      • STAR magnet ramping/chiller problems

      • Early dump, YS3 trip/bad rebucketing

      • QLI at dump, Yellow debunched beam

    • 1 extended store

      • Insufficient Tandem beam

  • Some YL cooling issues, off for last 3 stores

Outstanding items

  • Yellow, Blue store conditions

  • Blue injection efficiency

  • Storage cavity voltage calibration

    • 1 test ramp

  • BV cooling pickup

    • Maintenance day item

  • LLRF firmware upgrade

    • 4 hrs, week of 6/13

  • AGS ag17-qgt magnet failure

    • Deferred to shutdown


Next 48 hours 6 1 3

Next 48 hours 6/1-3

  • Complete maintenance

    • BV cooling pickup

    • YL cooling kicker motion control

    • YS2, YS3 RF

    • Yellow collimator pindiodes

  • Resume Physics

    • STAR, PHENIX polarity switches

    • Vernier scans Thursday or Friday: early in store

  • Machine Development Thursday:

    • End of store: test beta* squeeze, 2 IRs 4 planes

    • Starts between daytime stores

    • Store conditions

      • Working point scans (dedicated ramp)

      • Storage voltage calibration

  • Status meetings M-W-F


Heavy ion run update 5 23 25

Heavy Ion Run Update 5/23-25

  • APEX completed Wednesday

    • Attempt at fixing post-γt ramp loss

  • Physics running, 4 stores

    • 3 failed ramps

      • Early loss attributed to y-qtrim?

      • FEC reboot and subsequent steering error

      • Fast loss ~10 sec after γt

    • 1 aborted store

      • Unexplained permit drop 6a to 7a

    • AGS vacuum failure+Blue injection kicker

      • Other work completed concurrently

    • Machine Development deferred

    • Some YV cooling issues(limit switch fixed)

Outstanding items

  • AGS ag17-qgt magnet failure

    • Long maintenance day or shutdown repair

  • Yellow, Blue store conditions

  • Blue injection efficiency

  • Storage cavity voltage calibration

    • 1 test ramp

  • BV cooling pickup

    • Maintenance day item

  • LLRF firmware upgrade

    • 4 hrs, week of 6/13


Holiday weekend 5 27 31

Holiday weekend 5/27-31

  • Continue Physics

  • Machine Development items:

    • Store conditions

      • Working point scans (dedicated ramp)

      • Yellow negative chromaticity? Blue?

      • Beta* squeeze – prepare for automation.

    • More beam

    • Ramp Efficiency

      • Ramp quicker after fill

  • Status meetings M-W-F

  • Run Coordinator will be off-island but on cell...

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.


Heavy ion run update 5 23 251

Heavy Ion Run Update 5/23-25

  • Physics running, 7 stores

    • No aborted stores (2 early dumps, 1 late dump).

    • 5 failed ramps

      • Sextupole ramp rate (2x)

      • Post-transition loss (3x)

    • Accesses for repairs: YA1 cavity, stochastic cooling amplifiers, sextupole feedthrough ground current, collimator pin diodes…

    • QD controls problems (2x)

  • 3rd test of beta squeeze at store

    • 4 planes successfully squeezed at once

    • Requires more beam time to measure optics, interpret resulting rate changes.

  • Vernier scans conducted

  • Outstanding items

    • Blue, Yellow store conditions

      • MD Thursday

    • Storage cavity voltage calibration

      • Requires test ramp

    • Blue injection efficiency

      • Fine tuning

      • Offline evaluation

    • AGS: ag17-qgt, tune quad

    • ATR synchro errors

    • BV cooling pickup

      • Open vacuum to repair

    • LLRF firmware upgrade

      • 4 hrs week of 6/13


Next 48 hours 5 25 27

Next 48 hours 5/25-27

  • APEX completes 2359 tonight

  • Continue Physics

  • Machine Development Thursday

    • Store conditions

      • Working point scans (dedicated ramp)

      • Yellow negative chromaticity? Blue?

      • Beta* squeeze – prepare for automation.

    • More beam

    • Ramp Efficiency

      • Ramp quicker after fill

  • Status meetings M-W-F

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.


Heavy ion run update 5 20 23

Heavy Ion Run Update 5/20-23

  • Physics running, 14 stores

    • Only 2 failed ramps (Yellow abort prefire)

    • No aborted stores.

  • All stochastic cooling systems on

    • Still some kicker insertion issues

  • Additional test of beta squeeze at store (successful)

    • Requires more beam time to measure optics, interpret resulting rate changes.

  • Outstanding items

    • YA1 cavity problem

    • Source/injector intensity

    • Blue, Yellow store conditions

      • Ready for more beam?

    • Storage cavity voltage calibration

      • Requires test ramp

    • Blue injection efficiency

      • Fine tuning

      • Offline evaluation


Next 48 hours 5 23 25

Next 48 hours 5/23-25

  • Continue Physics

    • AGS access behind store for RF?

    • Store length with cooling?

    • Continue stochastic cooling development

  • Machine Development items for improved luminosity

    • More beam

    • Store conditions

      • Working point scans (dedicated ramp)

      • Yellow negative chromaticity? Blue?

      • Beta* squeeze – prepare for automation.

    • Ramp Efficiency

      • Ramp quicker after fill

  • Status meetings M-W-F

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.


Heavy ion run update 5 18 20

Heavy Ion Run Update 5/18-20

  • Maintenance completed

    • YS5, BS5 in operation

    • Booster B3 RF modified

  • Physics running, 8 stores

    • Storage cavity trip at rebucketing

    • Loss at rebucketing

    • AGS extraction issue, early dump

    • Power dip

    • Loss at transition

  • All stochastic cooling systems on but…

    • YL slip detection issue

    • BV broken switch

  • Beta squeeze at store tested

  • Outstanding items

    • Source/injector intensity

    • Tenuous Yellow store conditions

    • Storage cavity voltage calibration

    • Blue injection efficiency

      • Fine tuning

      • Offline evaluation


Rhic heavy ion run 11

YL, BL

YL, BL, YV, BV

YL, VL, YV, BV

YL, BL, YV


Rhic heavy ion run 11

Measured tune shift at rebucketing


Weekend outlook 5 20 23

Weekend outlook 5/20-23

  • Continue Physics

    • Store length with cooling?

  • Stochastic cooling

    • Goal: all 4 operational this weekend

  • Machine Development items for improved luminosity

    • Store conditions

      • Yellow negative chromaticity? Blue?

      • Check y5-lm3.2 loss, collimation setup

    • Ramp Efficiency

      • Ramp quicker after fill

    • More beam

  • Status meetings M-W-F

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.


Heavy ion run update 5 16 18

Heavy Ion Run Update 5/16-18

  • Physics running, 7 stores

    • Yellow losses after rebucketing (5x)

    • bo3-sxd1 trip, loss after rebucketing

    • ATR cogging problem during/after MD

    • Other LLRF issue?

      • “Cured” by hysteresis

  • Machine Development

    • Corrected Blue injection

  • Yellow vertical cooling operational, but losses at kicker exceed limits

  • Outstanding items

    • Blue injection efficiency

      • Fine tuning

      • Offline evaluation

    • Booster B3 RF, deferred troubleshooting

    • YS5, BS5 cavity conditioning

      • Maintenance item

    • Storage cavity voltage calibration


Rhic heavy ion run 11

Injection efficiencies


Next 48 hours 5 18 20

Next 48 hours 5/18-20

  • Finish Maintenance

  • Continue Physics

    • Store length with partial cooling?

  • Stochastic cooling setup, Blue vertical

    • Every possible store

  • Machine Development items for improved luminosity

    • Store conditions

      • Yellow negative chromaticity? Blue?

      • Have we drifted from optimal working point?

    • Ramp Efficiency

      • Ramp quicker after fill

    • More beam

  • Status meetings M-W-F

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.


Heavy ion run update weekend 5 13 16

Heavy Ion Run Update weekend 5/13-16

  • Physics running, 12 stores

    • Landau cavity configuration issue

    • Controls server failure + restoration issue

    • Transition abort

    • Yellow store collimation

  • Blue & Yellow longitudinal cooling operational

  • Outstanding items

    • Blue injection efficiency

    • Booster B3 RF, deferred troubleshooting

    • YS5, BS5 cavity conditioning


This week s outlook 5 16

This week’s outlook 5/16-?

  • Continue Physics

    • Store length w/o cooling should be 4 hours.

  • Stochastic cooling setup, vertical

    • Every possible store

  • Machine Development items for improved luminosity

    • Ramp Efficiency

      • Blue injection efficiency

      • Ramp quicker after fill

      • Yellow beam decay spikes at rebucketing.

      • Longitudinal growth after transition – gap volts? Low intensity bunch?

    • More beam

  • Next meetings M-W-F

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.


Heavy ion run update 5 12 13

Heavy Ion Run Update 5/12-13

  • Physics running, 5 stores

    • Peak luminosities approaching best stores from last run

  • Yellow longitudinal cooling operational

  • Blue longitudinal tested, not yet operational

  • Outstanding items

    • BS2 cavity trip

    • Blue injection efficiency

      • Better, but still variable

    • Booster B3 RF, deferred troubleshooting

    • YS5, BS5 cavity conditioning

      • More conditioning for YS4, BS4?


Rhic heavy ion run 11

Longitudinal cooling on/off


Weekend 5 13 16

Weekend 5/13-16

  • Continue Physics

    • Store length w/o cooling should be 4 hours.

  • Stochastic cooling setup

    • Every possible store

  • Machine Development items for improved luminosity

    • Ramp Efficiency

      • Ramp quicker after fill

      • Blue injection efficiency

      • Transition efficiency – chromaticity? Yellow dipole feedback?

      • Yellow beam decay spikes at rebucketing.

    • More beam

  • Next meeting Monday noon in LCR.

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.


Heavy ion run update 5 10 12

Heavy Ion Run Update 5/10-12

  • Physics running, 4 stores

  • Stochastic cooling setup continues

  • APEX

  • Maintenance

    • YA2 opened, repaired

    • YS4, BS4 cavities in operation

    • LLRF failure post-maintenance

  • Outstanding items

    • Blue injection efficiency

      • Better, but still variable

    • Booster B3 RF, deferred troubleshooting

    • Vernier scans

    • YS5, BS5 cavity conditioning

      • More conditioning for YS4, BS4?


Rhic heavy ion run 11

Stores since Tuesday


Rhic heavy ion run 11

Injection efficiency


Rhic heavy ion run 11

IR4 Vacuum


Next 24 hours 5 11 12

Next 24 hours 5/11-12

  • Continue Physics

    • Store length w/o cooling should be 4 hours.

  • Stochastic cooling setup

    • Every possible store

  • Machine Development items for improved luminosity

    • Ramp Efficiency

      • Ramp quicker after fill

      • Transition efficiency – chromaticity? Yellow dipole feedback?

  • More vernier scans?

  • Meeting tomorrow noon in LCR.

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.


Heavy ion run update 5 9 10

Heavy Ion Run Update 5/9-10

  • Access for RF troubleshooting, experiments

  • Physics running, more long stores

    • PS problems

      • Main magnet

    • Abort kicker prefire

    • Stochastic cooling trip

  • Stochastic cooling setup continues

  • Outstanding items

    • YA2 cavity

    • AGS extraction efficiency

    • Blue injection efficiency

    • Booster B3 RF, deferred troubleshooting

    • Vernier scans

    • YS4, YS5, BS4, BS5 cavity conditioning


Next 24 hours 5 10 11

Next 24 hours 5/10-11

  • Continue Physics

    • Reduce store length w/o cooling?

  • Stochastic cooling setup

    • Every possible store

  • Machine Development items for improved luminosity

    • AGS extraction efficiency work (parasitic)

      • Still ~7% low?

    • Ramp Efficiency

      • Ramp quicker after fill

      • Transition efficiency – chromaticity?

  • Schedule vernier scans

  • APEX 0001-0800

  • Maintenance 0800-1600

  • NO MEETING TOMORROW

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.


Heavy ion run update weekend 5 6 9

Heavy Ion Run Updateweekend 5/6-9

  • Physics running but with few normal dumps

    • PS problems

      • Sextupole trip

      • Main magnet

      • UPS

    • Periodic orbit correction/lead flow

    • Abort kicker prefire

    • Stochastic cooling trip

    • YA2 RF problems

  • Stochastic cooling setup continues

  • Outstanding items

    • YA2 cavity

    • AGS extraction efficiency

    • Booster B3 RF, deferred troubleshooting

    • Vernier scans

    • YS4, YS5, BS4, BS5 cavity conditioning

    • WCM calibration?


Rhic heavy ion run 11

Weekend stores


Rhic heavy ion run 11

Rates and backgrounds


Rhic heavy ion run 11

WCM and DCCT


Next 24 hours 5 9 10

Next 24 hours 5/9-10

  • Troubleshoot YA2

  • Continue Physics

    • Reduce store length w/o cooling?

  • Stochastic cooling setup

    • Every possible store

  • Machine Development items for improved luminosity

    • AGS extraction efficiency work (parasitic)

    • Ramp Efficiency

      • Ramp quicker after fill

      • Transition efficiency – chromaticity?

  • Schedule vernier scans

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.


Heavy ion run update 5 5 6

Heavy Ion Run Update 5/5-6

  • Rebucketing operational

  • 2 stores overnight

  • Outstanding items

    • Ramp efficiency

      • 5-6% snapback loss

      • 3-8% transition loss

    • Storage cavity conditioning

      • New cavities not yet up to voltage

    • Booster B3 RF cavity

      • May limit our highest injector intensities

    • BS3 cavity

      • investigating


Weekend 5 6 9

Weekend 5/6-9

  • Ramp tuning

    • Transition losses – chromaticity?

    • Orbit after transition

    • Injection debunching – check cavity phasing?

    • Continue ramp collimation setup

  • Physics

    • Store length 4 hours?

  • Store setup

    • Stochastic cooling setup, longitudinal first

    • Check & adjust store chromaticity

    • Increase bunch intensity/number up to 1.3/111

  • Other items

    • Check injection drift correction

    • Optics measurements, Blue

    • Ramp chromaticity with Booster off?

  • WEEKEND MEETINGS -- cancelled


Heavy ion run update 5 4 5

Heavy Ion Run Update 5/4-5

  • Beam resumed ~2000 after maintenance

    • Bunch intensities approaching 1.3e9

  • YA1 cavity operational

  • Landau cavities on

  • Rebucketing setup started

    • Storage cavity phasing

  • Yellow store chromaticity measured

    • Suggests negative horizontal

  • Yellow tune scan at store

    • Best Yellow: (.2319, .2071)

    • Blue: (.2310, .2147)

  • b2-dhx, sector 5 qgt problems

  • Outstanding items

    • Ramp efficiency

      • 5-6% snapback loss

      • 3-8% transition loss

    • Storage cavity conditioning

      • New cavities not yet up to voltage

    • WCM rebucket error

    • Booster B3 RF cavity

      • May limit our highest injector intensities


Rhic heavy ion run 11

Ramp Efficiency


Rhic heavy ion run 11

Present working point


Next 24 hours 5 5 6

Next 24 hours 5/5-6

  • Store setup

    • Rebucket setup has highest priority

    • Stochastic cooling setup, longitudinal first

    • Check & adjust chromaticity

    • Increase bunch intensity/number up to 1.3/111

  • Overnight stores for experiments

    • Pending rebucket work

  • Other items

    • Continue ramp collimation setup

    • Check injection drift correction

    • Optics measurements

    • Ramp chromaticity with Booster off?


Heavy ion run update 5 3 4

Heavy Ion Run Update 5/3-4

  • Collisions established

    • 79x79, 1e11

  • Initial collimation complete

  • 10 Hz feedback running

  • Gap cleaning operational

  • Outstanding items

    • YA1 cavity

      • Repair made, opened vacuum during maintenance today

    • Misc elog issues

      • Ramp efficiency entry

      • RatView scales

    • Booster B3 RF cavity


Rhic heavy ion run 11

Store tunes


Rhic heavy ion run 11

IPM Emittance


Next 24 hours 5 4 5

Next 24 hours 5/4-5

  • Finish maintenance, restore beam

    • Evaluate YA1 repair

  • Store setup

    • Rebucket setup

    • Stochastic cooling setup, longitudinal first

    • Increase bunch intensity/number up to 1.3/111

  • Other items

    • Landau cavities

    • Check injection drift correction

    • Continue ramp collimation setup

    • Store chromaticity, Yellow

    • Optics measurements

    • Ramp chromaticity with Booster off?


Heavy ion run update 5 2 3

Heavy Ion Run Update 5/2-3

  • Beam to store on first ramp

  • Ramp development finishes major work in ~3 hours

  • Completed items:

    • Gamma jump, yellow dipole feedback

    • BPM timing @ injection, store

    • Tune/coupling/orbit feedback employed

    • Tune/coupling replay exercised

    • Ramp chromaticity measurement

  • Outstanding items

    • YA1 cavity

      • Presently focusing on external issues

    • Note change in global beam sync timing

      • 26 ns Blue

      • 24 ns Yellow

    • WCM transition

    • Misc elog issues

      • Ramp efficiency entry

      • Injection statistics?

      • RatView scales


Rhic heavy ion run 11

First ramp attempt


Rhic heavy ion run 11

Blue ramp chromaticity


Rhic heavy ion run 11

Yellow ramp chromaticity


Next 24 hours 5 3 4

Next 24 hours 5/3-4

  • Ramp development, store setup until 0700 Wednesday

    • Investigate YA1 issue (access if necessary)

    • Increase bunch intensity/number up to 56

    • Ramp collimation

    • RhicInjection orbit feedback test

    • RF cogging on ramp

    • Separation bump removal on ramp

    • Steer for collisions

    • Rebucket? Stochastic cooling?

    • Ramp chromaticity with Booster off?


Heavy ion run update 4 29 5 2

Heavy Ion Run Update 4/29-5/2

  • Low energy run completed at 0800 this morning

  • Weekend troubles:

    • Lead flow interlock

    • AGS L10 RF cavity

  • Outstanding items

    • Injection efficiency

      • Lower than normal injection

      • Limits injected intensity

    • Booster B3 RF, deferred troubleshooting


Next 24 hours 5 2 3

Next 24 hours 5/2-3

  • Injection beam development 1100-1300

  • PS ramp testing, Booster B3 RF troubleshooting ~1300-1500 Monday.

  • Setup with beam resumes 1500 today.

  • AP shifts resume

    • Re-establish injection

    • Instrumentation check

    • Ramp Development


Heavy ion run update 4 28 29 11

Heavy Ion Run Update 4/28-29/11

  • PHENIX zero field run completed, settings reverted.

  • Second set of vernier scans conducted.

  • Power dip 1550-0030

    • Booster, AGS ring accesses required

  • Outstanding items

    • Injection efficiency

      • Lower than normal injection

      • Limits injected intensity

    • Booster B3 RF, deferred troubleshooting

    • Vernier scans


Weekend 4 29 5 2

Weekend 4/29-5/2

  • Continue Physics

    • Store length = 30 min

  • Last vernier scans

  • APEX 0800-1400 Monday

  • √sNN=200 GeV setup begins 1400 Monday

    • Ramp Au11v3::injection→Au11v1::injection with beam, feedback

  • PS ramp testing, Booster B3 RF troubleshooting ~1600 Monday.

  • AP shifts resume Monday evening


Heavy ion run update 4 27 28 11

Heavy Ion Run Update 4/27-28/11

  • PHENIX zero field run under way.

  • Outstanding items

    • Injection efficiency

      • Lower than normal injection

      • Limits injected intensity

    • Booster B3 RF, deferred troubleshooting

    • Vernier scans


Rhic heavy ion run 11

Run-10 Au-Au


Rhic heavy ion run 11

Run-11 Au-Au


Next 24 hours 4 28 29

Next 24 hours 4/28-29

  • Resume normal stores

  • Continue Physics

    • Store length = 30 min

  • More vernier scans

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.


Heavy ion run update 4 26 27 11

Heavy Ion Run Update 4/26-27/11

  • Initial vernier scans completed

    • Optimized PHENIX, repositioned STAR

  • Booster beamstop failure 2130-0600

    • Replaced bo6-qs3, yi6-sx3

    • Repaired bi8-qf3

    • Sublimated sector 4 vacuum

  • Outstanding items

    • Injection efficiency

      • Limits injected intensity

    • Orbit control

      • High RMS since hysteresis

    • Booster B3 RF, deferred troubleshooting

    • Vernier scans


Next 24 hours 4 27 28

Next 24 hours 4/27-28

  • Examine closed orbit, injection efficiency

  • Continue Physics

    • Store length = 30 min

  • More vernier scans

  • PHENIX zero field run

    • Prior to tour access

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.

  • NO MEETING TOMORROW


Heavy ion run update 4 25 26 11

Heavy Ion Run Update 4/25-26/11

  • MD for Yellow

    • Delayed by Booster failure

    • Yellow orbit work: 1 hour

      • Injection efficiency up > 10%

      • No BLM issues since

    • Initial collimation work: 1 hour But…

  • Blue orbit shifted

    • Why? No evidence found yet

    • Corrected blue steering for collisions

    • Re-collimated.

    • Eventually, corrected closed orbit again and resteered, collimated.

  • Outstanding items

    • Injection efficiency

      • Limits injected intensity

    • Orbit control

      • Can we correct orbit drift? Should we?

      • BPM issues?

    • Booster B3 RF, deferred troubleshooting

    • Vernier scans

    • bo6-qs3 failure


Next 24 hours 4 26 27

Next 24 hours 4/26-27

  • Continue Physics

    • Store length = 30 min

  • Schedule vernier scans

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.


Heavy ion run update weekend 4 22 25 11

Heavy Ion Run Update Weekend 4/22-25/11

  • Mostly continuous running

    • Over 80 stores since Friday afternoon

    • Store length reduced to 30 min

      • Optimal ~15-20 min

  • Scanned RF voltages

  • PHENIX “vertex shift”

    • Determined to be collimator backsplash from opposite ring

  • Outstanding items

    • Beam lifetime

    • Injection efficiency

      • Limits injected intensity

    • Orbit control

    • Booster B3 RF, deferred troubleshooting

    • Vernier scans

    • bo6-qs3 failure

    • Instrumentation setup

      • IPM, WCM


Next 24 hours 4 25 26

Next 24 hours 4/25-26

  • Continue Physics

    • Store length = 30 min

  • Machine Development items for improved luminosity

    • Work on orbit feedback management

      • Yellow bend trim adjustment

    • Injection efficiency work

    • Beam lifetime work

      • Change working point

      • Better RF voltage choice (parasitic)?

      • Octupoles (parasitic)?

    • 10 Hz GOFB (parasitic)

  • Schedule vernier scans

  • AP shifts: on call

    • ...but be prepared to log in remotely or come in if necessary.


Heavy ion run update fri 4 22 11

Heavy Ion Run Update Fri 4/22/11

  • Booster B3 RF cavity running, issues unresolved

    • Will be of concern when we start to push for maximum injector performance

  • Beam restored, captured in both rings

  • Tuned up to nominal intensities

  • Initial collision steering and collimation successful

  • Outstanding items

    • Bunched beam lifetime

    • Booster B3 RF, deferred troubleshooting

    • Store automation


Weekend 4 22 24

Weekend 4/22-24

  • Start Physics

    • Store length = 1 hour?

  • Improve store conditions, store turnaround times.

    • Store automation

    • Optimize rates, minimize backgrounds

    • Increase intensities

  • AP shifts: Friday in MCR, Saturday/Sunday on call

    • ...but be prepared to log in remotely or come in if necessary.

  • NO meeting Sunday. Saturday?


Heavy ion run update thu 4 21 11

Heavy Ion Run Update Thu 4/21/11

  • 28 MHz RF repairs, q7 polarity reversal complete Wed ~2300

  • Beam restored to both rings

  • Intensities too low for RF capture

  • Beam from injectors unstable

  • Node card/UPS failure, 1007C

  • Outstanding items

    • Tunes far from design

    • Corrector strength/avg offset too high?

    • Booster B3 RF, AGS bunch merge


Rhic operations

RHIC operations


Injector performance

Injector performance


Atr transport

ATR transport


Next 24 hours 4 21 22

Next 24 hours 4/21-22

  • Repair Booster B3 RF cavity

  • PHENIX access for tour

  • Resume RHIC injection setup

    • Circulating beam

      • Check set/measured tunes vs. Au104/6/7

      • Revert orbits to pp or med/hi Au ramps?

    • RF capture

    • Lifetime/intensity tuning


Heavy ion run update wed 4 20 11

Heavy Ion Run Update Wed 4/20/11

  • 28 MHz RF acceleration cavity mechanical tuner problem

    • fRF at 9 GeV: 27.9996 MHz

    • fRF at 9.8 GeV normal injection: 28.0232 MHz

    • Cavity lower limit: 28.070 MHz

  • Setup items completed:

    • AGS-RHIC synchro

    • AGS extraction, ATR transport (twice)

    • 9.8 GeV lattice loaded

    • RHIC first turn

    • Injection kicker setup, circulating beam Blue and Yellow

    • Blue RF capture


Next 24 hours 4 20 21

Next 24 hours 4/20-21

  • Access to repair RF cavity tuners

  • Reverse q7 polarities, switch to IBS lattice.

  • Address other maintenance items

    • AGS main magnet motor maintenance

  • Resume RHIC injection setup

    • 9.8 GeV, Au11v3

    • First turn

    • Circulating beam

    • RF capture

    • Lifetime/intensity tuning


  • Login